A 31-inch-square TV is on sale at the local electronics store. If 31 inches measure of the diagonal of the screen, use the Pythagorem to find the legnth of the side of the screen.
Found 2 solutions by richwmiller, edjones:
Answer by richwmiller(17219)

a^2+b^2=c^2
where a=b and c=31
2a^2=31^2
2a^2=961
a=31/sqrt(2)=21.9203

Answer by edjones(8007)

Let a=one of the sides
2a^2=31^2 Pythagoras
a^2=961/2=480.5
a=21.92"
.
Ed